# IQ-Aware Precoding for Atomic MIMO Receivers

### v1

@@ V-B1 Problem statement
We then investigate the SC-PS network. In this architecture, each RF is connected to a subset of the BS array with $K=N_{t}/N_{\rm RF}$ antennas. The number of required PSs is thus $N_{t}$. With much reduced PSs, the SC architecture exhibits lower spectral efficiency than the FC one, but is more energy-efficient and easier to implement in practice.
To simplify notation, we employ the same expressions for the digital precoder $\bar{{\mathbf{D}}}$ and analog precoder $\bar{{\mathbf{A}}}$ as used in Section V-A. The major difference lies in that ${\mathbf{A}}={\mathbf{A}}_{I}+j{\mathbf{A}}_{Q}$ is a block diagonal matrix in the SC-PS structure:
@@
where ${\mathbf{p}}_{n}=[{\rm exp}(j\theta_{n,1}),{\rm exp}(j\theta_{n,2}),\cdots,{\rm exp}(j\theta_{n,K})]^{T}$ represent the PSs connected to the $n$-th RF chain. One can verify that ${\mathbf{A}}^{H}{\mathbf{A}}=K{\mathbf{I}}_{N_{\rm RF}}$ and accordingly $\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}=K{\mathbf{I}}_{2N_{\rm RF}}$ without difficulty. Bearing this property, the total power constraint is simplified to $\frac{1}{2}\mathsf{Tr}\{\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}\bar{{\mathbf{D}}}^{T}\bar{{\mathbf{A}}}^{T}\}=\frac{K}{2}\mathsf{Tr}\{\bar{{\mathbf{D}}}\bar{{\mathbf{D}}}^{T}\}=P$. Henceforth, the precoding design problem is formulated as
@@
$(\mathcal{P}6)\quad\mathrm{s.t.}\>\>$ ${\mathbf{A}}={\rm blkdiag}\{{\mathbf{p}}_{1},{\mathbf{p}}_{2},\cdots,{\mathbf{p}}_{N_{\rm RF}}\},$ (71a)
@@
Notably, the total power constraint is now independent of the analog precoder $\bar{{\mathbf{A}}}$. This fact enables us to directly utilize the alternating minimization approach to solve problem $(\mathcal{P}6)$.
@@ V-B2 Algorithm design
Similar to the FC algorithm, our SC hybrid precoding algorithm alternatively updates the analog and digital precoders while fixing the other one. The optimization details are presented below.
Algorithm 2 : IQ-aware hybrid precoding with SC-PS network
@@
i) Fix $\bar{{\mathbf{D}}}$ and optimize $\bar{{\mathbf{A}}}$: Given $\bar{{\mathbf{D}}}$, we leverage matrix techniques to derive the optimal $\bar{{\mathbf{A}}}$ that minimizes (71). By substituting $\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}=K{\mathbf{I}}_{2N_{\rm RF}}$ into (71), the objective function is rewritten as
@@
where $C_{3}=\mathsf{Tr}\{\bar{{\mathbf{F}}}^{T}\bar{{\mathbf{F}}}\}+K\mathsf{Tr}\{\bar{{\mathbf{D}}}^{T}\bar{{\mathbf{D}}}\}$. To express (72) as a function of ${\mathbf{A}}_{I}$ and ${\mathbf{A}}_{Q}$ explicitly, we notice that the matrix $\bar{{\mathbf{A}}}$ has an equivalent expression:
$\bar{{\mathbf{A}}}$ $=\underbrace{\left(\begin{array}[]{cccc}{\mathbf{I}}_{N_{t}}&{\mathbf{0}}&{\mathbf{0}}&-{\mathbf{I}}_{N_{t}}\\ {\mathbf{0}}&{\mathbf{I}}_{N_{t}}&{\mathbf{I}}_{N_{t}}&{\mathbf{0}}\\ \end{array}\right)}_{{\mathbf{U}}}\left(\begin{array}[]{cc}{\mathbf{A}}_{I}&{\mathbf{0}}\\ {\mathbf{A}}_{Q}&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{A}}_{I}\\ {\mathbf{0}}&{\mathbf{A}}_{Q}\\ \end{array}\right)$ (79)
$={\mathbf{U}}\left(\begin{array}[]{cc}{\mathbf{A}}_{e}&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{A}}_{e}\\ \end{array}\right),$ (82)
where ${\mathbf{A}}_{e}\overset{\Delta}{=}[{\mathbf{A}}_{I}^{T},{\mathbf{A}}_{Q}^{T}]^{T}$. Furthermore, let the block matrix components of ${\mathbf{U}}^{T}\bar{{\mathbf{F}}}\bar{{\mathbf{D}}}^{T}$ be ${\mathbf{Y}}_{k\ell}\in\mathbb{R}^{2N_{t}\times N_{\rm RF}},k\in\{1,2\},\>\ell\in\{1,2\}$. We can rewrite $\mathsf{Tr}\{\bar{{\mathbf{D}}}\bar{{\mathbf{F}}}^{T}\bar{{\mathbf{A}}}\}$ in (72) as
$\mathsf{Tr}\{\bar{{\mathbf{D}}}\bar{{\mathbf{F}}}^{T}\bar{{\mathbf{A}}}\}=\mathsf{Tr}\left\{\left(\begin{array}[]{cc}{\mathbf{Y}}_{11}^{T}&{\mathbf{Y}}_{21}^{T}\\ {\mathbf{Y}}_{12}^{T}&{\mathbf{Y}}_{22}^{T}\end{array}\right)\left(\begin{array}[]{cc}{\mathbf{A}}_{e}&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{A}}_{e}\\ \end{array}\right)\right\}$ (87)
$=\mathsf{Tr}\{({\mathbf{Y}}_{11}+{\mathbf{Y}}_{22})^{T}{\mathbf{A}}_{e}\}\overset{(a)}{=}\mathsf{Tr}\left\{({\mathbf{Y}}_{I},{\mathbf{Y}}_{Q})^{T}\left(\begin{array}[]{c}{\mathbf{A}}_{I}\\ {\mathbf{A}}_{Q}\end{array}\right)\right\}$ (90)
@@
This completes the update of ${\mathbf{A}}$.
ii) Fix $\bar{{\mathbf{A}}}$ and optimize $\bar{{\mathbf{D}}}$: For a fixed analog precoder, the optimization of $\bar{{\mathbf{D}}}$ is a quadratic programming problem with quadratic constraint. By applying the Karush-Kuhn-Tucker (KKT) conditions, the optimal solution is derived as
$\bar{{\mathbf{D}}}^{\star}=\left(\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}+\lambda^{\star}{\mathbf{I}}_{2N_{\rm RF}}\right)^{-1}\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{F}}}=\frac{1}{K+\lambda^{\star}}\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{F}}},$ (93)
wherein the Lagrange multiplier $\lambda^{\star}$ is selected to fulfill the total power constraint. Accordingly, substituting (93) into (71b), we arrive at
$\bar{{\mathbf{D}}}^{\star}=\sqrt{\frac{2P}{K\|\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{F}}}\|_{F}^{2}}}\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{F}}}.$ (94)
As a result, a step-by-step summary of the IQ-aware hybrid precoding algorithm for SC-PS structure is provided in Algorithm 2.
@@ V-C1 Convergence analysis
For both Algorithms 1 and 2, the updated solutions in each iteration are all optimal to their respective sub-problems. Thereafter, our algorithms can monotonically decrease the objective functions $\|\widetilde{{\mathbf{F}}}\widetilde{{\mathbf{D}}}^{T}-\bar{{\mathbf{A}}}\|_{F}^{2}$ and $\|\bar{{\mathbf{F}}}-\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}\|_{F}^{2}$. Moreover, these two functions have a lower bound of zero. By invoking the monotone convergence theorem, the convergence of Algorithms 1 and 2 are guaranteed.
@@ V-C2 Computational complexity analysis
To evaluate the computational complexities of Algorithms 1 and 2, we access the number of scalar multiplications required by each algorithm.
To begin with, consider one iteration of Algorithm 1. Updating ${\mathbf{A}}_{I}$ and ${\mathbf{A}}_{Q}$ necessitates the calculation of matrices ${\mathbf{Z}}_{I}$ and ${\mathbf{Z}}_{Q}$, which has a complexity of $\mathcal{O}(N_{t}N_{\rm RF}^{2})$. Next, the optimization of $\widetilde{{\mathbf{D}}}_{u}$ and $\widetilde{{\mathbf{F}}}_{c}$ involves the calculation of matrix $\widetilde{{\mathbf{F}}}^{T}\widetilde{{\mathbf{A}}}$ and its SVD, resulting in a complexity on the order of $\mathcal{O}(N_{t}N_{\rm RF}^{2}+N_{\rm RF}^{3})$. In summary, the overall complexity of Algorithm 1 is $\mathcal{O}(I_{0}(N_{t}N_{\rm RF}^{2}+N_{\rm RF}^{3}))$, where $I_{0}$ is the number of iterations required for convergence.
In addition, consider one iteration of Algorithm 2. The complexity for updating ${\mathbf{A}}_{I}$ and ${\mathbf{A}}_{Q}$ primarily results from computing the expression of matrix ${\mathbf{Y}}$. This step has a complexity of $\mathcal{O}(N_{t}N_{\rm RF}N_{s})$. Subsequently, the complexity of updating $\bar{{\mathbf{D}}}$ by (93) is on the order of $\mathcal{O}(N_{t}N_{\rm RF}N_{s}+N_{\rm RF}N_{s}^{2})$. To summarize, the computational complexity of Algorithm 2 is $\mathcal{O}(I_{0}(N_{t}N_{\rm RF}N_{s}+N_{\rm RF}N_{s}^{2}))$.
@@ VI-A Simulation Settings
The default simulation settings are as follows. The Rydberg energy levels are set as $62D_{5/2}$ and $64P_{3/2}$ to measure $\omega=2\pi\times 27.7\>{\rm GHz}$ radio frequency signals. The wavelength is $\lambda=\frac{2\pi c}{\omega}=10.83\>{\rm mm}$. Utilizing the Python package [36], the transition dipole moment $\bm{\mu}_{eg}$ over the states $62D_{5/2}$ and $64P_{3/2}$ is calculated as $[0,789.107\>qa_{0},0]^{T}$, wherein $a_{0}=5.292\times 10^{-11}$ m specifies the Bohr radius and $q=1.602\times 10^{-19}\>{\rm C}$ is the charge of an electron. The atomic antennas are deployed as a uniform linear array (ULA) with an equal antenna spacing $d=10\>{\rm mm}$. The transmitter-to-receiver channel coefficients are generated using the standard Saleh-Valenzuela multi-path channel model. Specifically, the number of paths is set to $L=10$. For each path, the complex coefficient $\rho_{mnl}e^{j\varphi_{mnl}}$ is generated by the equation $\rho_{mnl}e^{j\varphi_{mnl}}=\alpha_{l}e^{j\frac{md\sin\vartheta_{l}}{\lambda}}$ according to the ULA geometry. Here, the path gain $\alpha_{l}$ is sampled from the complex Gaussian distribution $\mathcal{CN}(0,1)$ and the incident angle $\vartheta_{l}$ is sampled from the uniform distribution $\mathcal{CN}(-90^{\circ},90^{\circ})$. Besides, the polarization direction $\bm{\epsilon}_{nkl}$ is randomly sampled on a unit circle perpendicular to the incident angle. As for the LO-to-receiver channel, since the LO-to-receiver distance is quite small, we fix the path losses $\{\rho_{R,m}\}$ as $\rho_{R}$ and generate the phases $\{\varphi_{R,m}\}$ from the distribution $\mathcal{U}(0,2\pi)$. The SNR is defined as $\frac{P}{\sigma^{2}}$, while the receive SNR is defined as
@@
VI-B DoF of Atomic Receivers
To begin with, the DoF of atomic MIMO receivers under high SNR regimes is evaluated in Fig. 5. We conduct 1000 Monte-Carlo trails to simulate the derivative of the ergodic capacity $C_{e}$ w.r.t. $\log{\rm SNR}$ ). The plotted curves include 1) the simulated DoF of a classic receiver when the phase of ${\mathbf{H}}{\mathbf{x}}+{\mathbf{r}}+{\mathbf{w}}$ is known, whose DoF is theoretically $\min\{N_{t},N_{r}\}$; 2) the simulated DoF of an atomic receiver with a theoretical DoF of $\min\{N_{t}/2,N_{r}\}$; and 3) the simulated DoF of in-phase transmission, where only the in-phase symbols, ${\mathbf{s}}_{I}$, are transmitted and its theoretical DoF is $\min\{N_{t}/2,N_{r}/2\}$. The number of the transmit antennas is fixed as $N_{t}=2$, while the number of receive antennas, $N_{r}$, grows from $2$ to $4$. We can observe from Fig. 5 that, with the increment of $N_{r}$, the DoF of an atomic receiver gradually climbs from 1 to 2. Thus, its capacity limit resembles the in-phase transmission when $N_{r}=N_{t}=2$ and becomes equivalent to the classic receiver when $N_{r}=2N_{t}=4$. This observation is consistent with our analytical results in Section III. Even if an atomic receiver is hard to measure the phase of incident EM waves, it experiences no loss in DoF with the help of a strong reference signal and sufficient number of receive antennas. By further considering that the atomic receiver has a stronger sensitivity and thus has a much higher receive SNR, it has the potential to outperform the classic counterpart.

### v2

@@ V-B1 Problem statement
To show that our proposed IQ-aware hybrid precoding is feasible to general analog circuits, we now turn to the investigation of SC-PS networks. In this architecture, each RF is connected to a subset of the BS array with $K=N_{t}/N_{\rm RF}$ antennas. The number of required PSs is thus $N_{t}$. With much reduced PSs, the SC architecture exhibits lower spectral efficiency than the FC one, but is more energy-efficient. The sub-connected feature states that the analog precoder ${\mathbf{A}}={\mathbf{A}}_{I}+j{\mathbf{A}}_{Q}$ is now a block diagonal matrix:
@@
where ${\mathbf{p}}_{n}=[{\rm exp}(j\theta_{n,1}),\cdots,{\rm exp}(j\theta_{n,K})]^{T}$ represent the PSs connected to the $n$-th RF chain. Thanks to this block diagonal property, one can verify that ${\mathbf{A}}^{H}{\mathbf{A}}=K{\mathbf{I}}_{N_{\rm RF}}$ and $\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}=K{\mathbf{I}}_{2N_{\rm RF}}$ with ease. These facts automatically releases $\bar{{\mathbf{A}}}$ from the total power constraint because $\frac{1}{2}\mathsf{Tr}\{\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}\bar{{\mathbf{D}}}^{T}\bar{{\mathbf{A}}}^{T}\}=\frac{K}{2}\mathsf{Tr}\{\bar{{\mathbf{D}}}\bar{{\mathbf{D}}}^{T}\}=P$. Henceforth, the SC hybrid precoding design problem is formulated as
@@
$\mathrm{s.t.}\>\>$ ${\mathbf{A}}={\rm blkdiag}\{{\mathbf{p}}_{1},{\mathbf{p}}_{2},\cdots,{\mathbf{p}}_{N_{\rm RF}}\},$ (66a)
@@
Clearly, problem (66) differs from the FC hybrid precoding problem (43) by two aspects: 1) ${\mathbf{A}}$ is block diagonal and 2) ${\mathbf{A}}$ is independent of the total power constraint. These differences enable us to alternatively optimize precoders $\bar{{\mathbf{A}}}$ and $\bar{{\mathbf{D}}}$ to solve (66), without the need of performing problem transformation as in Section V-A2.
@@ V-B2 Algorithm design
The step-by-step procedures of our SC hybrid precoding algorithm is summarized in Algorithm 2, while the details are explained below.
Algorithm 2 : IQ-aware hybrid precoding with a SC-PS network
@@
i) Fix $\bar{{\mathbf{D}}}$ and optimize $\bar{{\mathbf{A}}}$: To find the optimal $\bar{{\mathbf{A}}}$, we first substitute $\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}=K{\mathbf{I}}_{2N_{\rm RF}}$ into (66) and rewrite the objective function as
@@
where $C_{3}=\mathsf{Tr}\{\bar{{\mathbf{F}}}^{T}\bar{{\mathbf{F}}}\}+K\mathsf{Tr}\{\bar{{\mathbf{D}}}^{T}\bar{{\mathbf{D}}}\}$. Then, by defining ${\mathbf{Y}}_{k\ell}\in\mathbb{R}^{N_{t}\times N_{\rm RF}}$, $k\in\{1,2\},\>\ell\in\{1,2\}$ as the block matrix components of $\bar{{\mathbf{F}}}\bar{{\mathbf{D}}}^{T}$ such that
$\bar{{\mathbf{F}}}\bar{{\mathbf{D}}}^{T}={\left(\begin{array}[]{cc}{\mathbf{Y}}_{11}&{\mathbf{Y}}_{12}\\ {\mathbf{Y}}_{21}&{\mathbf{Y}}_{22}\end{array}\right)},$ (70)
we can expand $\mathsf{Tr}\{\bar{{\mathbf{D}}}\bar{{\mathbf{F}}}^{T}\bar{{\mathbf{A}}}\}$ as
$\mathsf{Tr}\left\{\left(\begin{array}[]{cc}{\mathbf{Y}}_{11}^{T}&{\mathbf{Y}}_{21}^{T}\\ {\mathbf{Y}}_{12}^{T}&{\mathbf{Y}}_{22}^{T}\end{array}\right)\left(\begin{array}[]{cc}{\mathbf{A}}_{I}&-{\mathbf{A}}_{Q}\\ {\mathbf{A}}_{Q}&{\mathbf{A}}_{I}\\ \end{array}\right)\right\}$ (75)
$=\mathsf{Tr}\{({\mathbf{Y}}_{11}+{\mathbf{Y}}_{22})^{T}{\mathbf{A}}_{I}+({\mathbf{Y}}_{21}-{\mathbf{Y}}_{12})^{T}{\mathbf{A}}_{Q}\}$
@@
ii) Fix $\bar{{\mathbf{A}}}$ and optimize $\bar{{\mathbf{D}}}$: For a fixed analog precoder, the optimization of $\bar{{\mathbf{D}}}$ is a quadratic programming problem with quadratic constraint. The optimal solution can be derived using the Karush-Kuhn-Tucker (KKT) conditions [40] as
$\bar{{\mathbf{D}}}^{\star}$ $=\left(\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{A}}}+\lambda^{\star}{\mathbf{I}}_{2N_{\rm RF}}\right)^{-1}\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{F}}}=\frac{1}{K+\lambda^{\star}}\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{F}}}$
=$\overset{(a)}{=}\sqrt{\frac{2P}{K\|\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{F}}}\|_{F}^{2}}}\bar{{\mathbf{A}}}^{T}\bar{{\mathbf{F}}}.$ (78)
wherein (a) holds by selecting a Lagrange multiplier $\lambda^{\star}$ to fulfill the total power constraint (66b). This completes the design of IQ-aware hybrid precoding with a SC-PS network.
@@ V-C1 Convergence analysis
For both Algorithms 1 and 2, the updated solutions in each iteration can monotonically decrease the objective functions $\|\widetilde{{\mathbf{F}}}\widetilde{{\mathbf{D}}}^{T}-\bar{{\mathbf{A}}}\|_{F}^{2}$ and $\|\bar{{\mathbf{F}}}-\bar{{\mathbf{A}}}\bar{{\mathbf{D}}}\|_{F}^{2}$. Moreover, these two functions have a lower bound of zero. By invoking the monotone convergence theorem, the convergence of Algorithms 1 and 2 are guaranteed.
@@ V-C2 Computational complexity analysis
To evaluate the computational complexities of Algorithms 1 and 2, we access the number of scalar multiplications required by each algorithm. Consider one iteration of Algorithm 1. Updating ${\mathbf{A}}_{I}$ and ${\mathbf{A}}_{Q}$ necessitates the calculation of matrices ${\mathbf{Z}}_{I}$ and ${\mathbf{Z}}_{Q}$, which has a complexity of $\mathcal{O}(N_{t}N_{\rm RF}^{2})$. Next, the optimization of $\widetilde{{\mathbf{D}}}_{u}$ and $\widetilde{{\mathbf{F}}}_{c}$ involves the calculation of matrix $\widetilde{{\mathbf{F}}}^{T}\widetilde{{\mathbf{A}}}$ and its SVD, resulting in a complexity on the order of $\mathcal{O}(N_{t}N_{\rm RF}^{2}+N_{\rm RF}^{3})$. In summary, the overall complexity of Algorithm 1 is $\mathcal{O}(I_{0}(N_{t}N_{\rm RF}^{2}+N_{\rm RF}^{3}))$, where $I_{0}$ is the number of iterations required for convergence. In addition, consider one iteration of Algorithm 2. The complexity for updating ${\mathbf{A}}_{I}$ and ${\mathbf{A}}_{Q}$ primarily results from computing the expression of matrix ${\mathbf{Y}}$. This step has a complexity of $\mathcal{O}(N_{t}N_{\rm RF}N_{s})$. Subsequently, the complexity of updating $\bar{{\mathbf{D}}}$ by (62) is on the order of $\mathcal{O}(N_{t}N_{\rm RF}N_{s}+N_{\rm RF}N_{s}^{2})$. To summarize, the computational complexity of Algorithm 2 is $\mathcal{O}(I_{0}(N_{t}N_{\rm RF}N_{s}+N_{\rm RF}N_{s}^{2}))$.
@@ VI-A Simulation Settings
The default simulation settings are as follows. The Rydberg energy levels are set as $62D_{5/2}$ and $64P_{3/2}$ to measure $\omega=2\pi\times 27.7\>{\rm GHz}$ radio frequency signals. The wavelength is $\lambda=\frac{2\pi c}{\omega}=10.83\>{\rm mm}$. Utilizing the Python package [41], the transition dipole moment $\bm{\mu}_{eg}$ over the states $62D_{5/2}$ and $64P_{3/2}$ is calculated as $[0,789.107\>qa_{0},0]^{T}$, where $a_{0}=5.292\times 10^{-11}$ m specifies the Bohr radius, and $q=1.602\times 10^{-19}\>{\rm C}$ is the charge of an electron. The uniform linear arrays (ULAs) are deployed at the transceiver with an equal antenna spacing $d=10\>{\rm mm}$. The transmitter-to-receiver channel coefficients are generated using the standard Saleh-Valenzuela multi-path channel model. Specifically, the number of paths is set to $L=10$. For each path, the complex coefficient $\rho_{mnl}e^{j\varphi_{mnl}}$ is generated by the equation $\rho_{mnl}e^{j\varphi_{mnl}}=\alpha_{l}e^{j\frac{d}{\lambda}(m\sin\vartheta_{l}+n\sin\varphi_{l})}$ according to the ULA geometry. Here, the path gain $\alpha_{l}$ is sampled from the complex Gaussian distribution $\mathcal{CN}(0,1)$. The incident and departure angles, $\vartheta_{l}$ and $\varphi_{l}$, are sampled from the uniform distribution $\mathcal{U}(-90^{\circ},90^{\circ})$. Besides, the polarization direction $\bm{\epsilon}_{nkl}$ is randomly sampled on a unit circle perpendicular to the incident angle. As for the LO-to-receiver channel, since the LO-to-receiver distance is quite small, we fix the path losses $\{\rho_{R,m}\}$ as $\rho_{R}$ and generate the phases $\{\varphi_{R,m}\}$ from the distribution $\mathcal{U}(0,2\pi)$. The SNR is defined as $\frac{P}{\sigma^{2}}$, while the receive SNR is defined as
@@
VI-B DoFs of Atomic Receivers
To begin with, the DoFs of atomic MIMO receivers under high SNR regimes are evaluated in Fig. 5. We conduct 1000 Monte-Carlo trails to simulate the derivative of the channel capacity $C$ w.r.t. $\log{\rm SNR}$ ). The plotted curves include: 1) the simulated DoFs of a classic receiver when the phase of ${\mathbf{H}}{\mathbf{x}}+{\mathbf{r}}+{\mathbf{w}}$ is known, whose DoFs are theoretically $\min(N_{t},N_{r})$; 2) the simulated DoFs of an atomic receiver with theoretical DoFs of $\min(N_{t}/2,N_{r})$; and 3) the simulated DoFs of in-phase transmission, where only the in-phase symbols, ${\mathbf{s}}_{I}$, are transmitted and its theoretical DoFs are $\min(N_{t}/2,N_{r}/2)$. The number of the transmit antennas is fixed as $N_{t}=2$, while the number of receive antennas, $N_{r}$, grows from $2$ to $4$. We can observe from Fig. 5 that, with the growth of $N_{r}$, the DoFs of an atomic receiver gradually increase from 1 to 2. Thus, its capacity limit resembles the in-phase transmission when $N_{r}=N_{t}=2$ and becomes equivalent to the classic receiver when $N_{r}=2N_{t}=4$. This observation is consistent with our analytical results in Section III. Even if an atomic receiver is hard to measure the phase of incident EM waves, it experiences no loss in DoF with the help of a strong reference signal and sufficient number of receive antennas. By further considering that the atomic receiver has a stronger sensitivity and thus has a much higher receive SNR, it has the potential to outperform the classic counterpart.